OpenStudy (anonymous):

Integral from 0 to 4*sqrt(2) of 1/sqrt(16-x^2). I'll rewrite below.

OpenStudy (anonymous):

\[\int\limits_{0}^{4\sqrt{2}}\frac{ dx }{ \sqrt{16-x^2} }\] I think I am trying to use the trig substitution of sin here.

OpenStudy (anonymous):

\[x=4\sin (\theta)\]

OpenStudy (raden):

let x = 4sinθ

OpenStudy (anonymous):

Is it \[(16 - 4\sin(\theta))^{1/2}\]

OpenStudy (anonymous):

Ahh wait. It's \[(16-4(\sin (\theta))^{2})^{1/2}\]

OpenStudy (raden):

x = 4sinθ dx = 4cosθ dθ x = 4sinθ then sqrt(16-16sin^2 θ)=sqrt(16(1-sin^2 θ))=sqrt(16cos^2 θ)= 4cosθ

OpenStudy (raden):

see your integral becomes